OpenStudy (anonymous):

evaluate the lim as x approaches infinity: sqrt(x^2+2x)-sqrt(x^2-x). wolframalpha is giving me 3/2 but i keep getting 1. (inf/inf). i know wolfram is right but i dont know why

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt{x^2+2x}+\sqrt{x^2-x}\]

OpenStudy (anonymous):

btw \[\frac{\infty}{\infty}\neq 1\] in general it is a statement about limits. if you get \[\frac{\infty}{\infty}\] i means you have to do more work

OpenStudy (anonymous):

oh alright i did not know that. so i did that first step and got 3x/(sqrt(x^2+2x)+sqrt(x^2-x))

OpenStudy (anonymous):

thats where i got inf/inf. so i dont know what to do from this point

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

best thing to do is reason it out \[\sqrt{x^2}=x\] so it is like having a denominator of \[2x\] giving you \[\frac{3x}{2x}=\frac{3}{2}\]

OpenStudy (anonymous):

if that is unsatisfactory, divide top and bottom by \[x\] to get \[\frac{3}{\sqrt{1+\frac{2}{x}}+\sqrt{1-\frac{1}{x}}}\] then you can see the 3/2

OpenStudy (anonymous):

ok awesome tyvm